I need help with this quadric formula
x2-2x+8=0

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

X^2-2X+8=0
(X-4)(X+2)=0
X-4=0
X=4 ANSWER.
X+2=0
X=-2 ANSWER.

Answer by edjones(8007)

x^2-2x+8=0
x^2-2x =-8
x^2-2x+1=-8+1 completing the square.
(x-1)^2=-7
x-1=+-sqrt(-7) sqrt of each side
x=1+-sqrt(7)i
.
Ed